Cajun Navy Boats, Inc. had the following financial information for 2018: Sales $756,000 Cost of Goods Sold $315,000 Operating Expenses $252,000 Operating Earnings $189,000 After-tax Operating Earnings $142,000 Current Liabilities $320,000 Minimum Acceptable Return 12% Average Cost of Capital 8% Average Invested Capital $1,890,000 Calculate the following items, showing your work: Return on Investment Capital Turnover Residual Income Economic Value Added
Answer #1

a)Return on investment = After-tax Operating Earnings / Average Invested Capital

                                    = 142000 / 1890000

                                    = .0751 or 7.51%

b)Capital turnover = Net sales /Average Invested Capital

                          = 756000/1890000

                          = .40

c)Residual income = After-tax Operating Earnings - (Minimum Acceptable Return *Asset invested)

                      = 142000- (.12*1890000)

                      = 142000 - 226800

                      = - 84800

d)Economic value added= After tax operating income -[weighted average cost of capital*(Total asset -current liabilities)]

         = 142000 - [.08*(1890000-320000)]

         =142000 -[.08*1570000]

          = 142000- 125600

          = 16400